# Robust Market Interventions

### v1

@@ 3. Recoverable structure
The first matrix, $\bm{D}_{\text{block}}$, is a block matrix that divides the $300$ products into $3$ equally-sized blocks. The entries of $\bm{D}_{\text{block}}$ are based on a $3\times 3$ matrix $\bm{C}$ that governs the pattern of interactions within blocks $(C_{ii}$) and across them ($C_{ij}$ with $i\neq j$). In particular, [^14] The symbol $\otimes$ denotes the Kronecker product.
@@
Here $\bm{J}$ is the matrix of ones. Products within each block are complements, and products across blocks are substitutes. For example, products in block 3 are highly substitutable with products in other blocks, whereas products in blocks 1 and 2 are only mildly substitutable with each other.
We can think of each block as a product category. In one category we have, for example, kitchen products, in another category we have digital entertainment products, and in the third category we have products related to sport activities. Within each of these categories, products are complements, but across the categories these products are substitutes. To take another example, we can think of three non-compatible operating systems that can be used to accomplish similar tasks; products within the same computer operating system are complements, and products across operating systems are substitutes.
The second term in eq. 7, a scaling of $-\bm{Z}\bm{Z}^{\prime}$, is a negative definite matrix that adds heterogeneity to the very regular pattern of interactions in $\bm{D}_{\text{block}}$. In particular, we take $\bm{Z}$ to be an $n\times 10$ matrix with rows drawn uniformly from the unit sphere. [^15] Note that entry $(i,i)$ of $\bm{Z}\bm{Z}^{\prime}$ is the norm of the $i$th row of $Z$, which, by construction, is $1$.
@@
where each entry of the error matrix $\bm{E}$—except for its diagonal entries, which are kept at zero—is drawn from $U[-1,1]$; this noise structure results in the upper bound $b(n)=n^{1/2}$ on $\|\bm{E}\|$.
We are interested in illustrating some basic statistical implications of $(M,\delta)$-recoverable structure that is strong enough relative to the noise $b(n)$—that is, with $M$ much larger than $b(n)$. In this example, the largest eigenvalue of $\bm{D}_{\text{block}}$ is much larger than $n^{1/2}$, and so when $\gamma$ is low, then the $\bm{D}$ of eq. 7 will also have some eigenvalues which are larger than $n^{1/2}$. In contrast, when $\gamma$ is high, then the main contributor to $\bm{D}$ is $-\bm{Z}\bm{Z}^{\prime}$ and, in this case, the largest eigenvalues of $\bm{D}$ turn out to be no larger than $n^{1/2}$. We now show that these two cases, low vs.  high $\gamma$, have very different implications for inference.
@@
First we consider a low value of $\gamma=0.3$. The two largest eigenvalues of $\bm{D}$, in absolute value, are approximately $130$ and $80$, and these are considerably larger than $b(n)=\sqrt{300}\approx 17$, whereas the third largest eigenvalue is roughly $1.2$; hence, $\mathcal{L}(\bm{D},M)$ with $M>b(n)$ equal to, e.g., $n^{2/3}$, is the subspace spanned by eigenvectors $\bm{u}^{1}$ and $\bm{u}^{2}$. As we will explain in Section 4.2, the Davis-Kahan Theorem tells us that we can use the noisy observation $\widehat{\bm{D}}$ to accurately estimate the eigenvectors in $\mathcal{L}(\bm{D},M)$. Here, for simplicity, we illustrate graphically how the first eigenvector of $\widehat{\bm{D}}$ is a good description of the first eigenvector of $\bm{D}$.
@@
In Section 5 we extend this example to illustrate our main result 1 stated in Section 4, and show how we use the condition that recoverable structure imposes on the quantity vector.
@@
4. Analysis of Robust Interventions
This section presents our main results, showing the existence of intervention rules that robustly achieve surplus improvements. To have a sense of what can be achieved at all, with or without noise, it is worth noting some basic properties of surplus under complete information:
Proposition 1.
The following hold true:
  1. (1)
    Under any intervention, surplus outcomes satisfy $\frac{1}{2}\dot{P}+\dot{C}=\dot{S}$.
  2. (2)
    For any intervention such that $\dot{C}\geq-\epsilon$, we have
    $\dot{P}+\dot{C}\leq(2+\epsilon)\dot{S}.$
Proof.
Consumer surplus satisfies the standard Marshallian formula $\dot{C}=-\bm{q}^{0}\cdot\dot{\bm{p}}$. Meanwhile, $\dot{P}=\bm{q}^{0}\cdot(\dot{\bm{p}}+\bm{\sigma})+(\bm{p}-\bm{c})\cdot\dot{\bm{q}}=2\bm{q}^{0}\cdot\dot{\bm{q}}$, where we have used the equilibrium condition $\bm{p}-\bm{c}={\bm{q}}$ from eq. 4. Finally, $\dot{S}=-{\bm{q}}^{0}\cdot\dot{\bm{c}}$ by definition. These facts together imply part (1), and part (2) is an immediate corollary. ∎
Part (1) implies that $\dot{P}+\dot{C}=2\dot{S}-\dot{C}$, and so the market’s surplus (the left-hand side) can potentially be increased without bound through the authority’s expenditures, a reduction in consumer surplus, or a combination.1616163 in Appendix B shows that the condition is not only necessary but also that, generically, all these outcomes can actually be achieved under complete information.
Part (2) focuses on the set of interventions subject to a lower bound on the surplus of consumers, and derives an upper bound on surplus gain achievable in that case—even by an omniscient authority. The constraint is natural for authorities that want to limit consumer welfare loss—e.g., if the authority is a platform that can lose customers, or a government that can lose political support.
Our main result, 1, examines what can be implemented robustly when the authority has partial information and the economy has a recoverable structure that is strong enough relative to the noise. It shows that there are interventions which robustly protect consumers from surplus loss and implement market surplus equal to the upper bound that an omniscient authority could achieve (the upper bound given by 1(2)).
Theorem 1.
Let $M(n)$ be an increasing sequence with $M(n)/b(n)\to_{n}\infty$, and fix $\delta>0$. Assume the set of market states has $(M(n),\delta)$-recoverable structure. For every $\epsilon>0$ and for every target expenditure $e>0$, the following properties can be simultaneously achieved $\epsilon$–robustly for sufficiently large $n$:
  1. (i)
    The sum of marginal consumer and producer surplus gains $\dot{P}+\dot{C}$ is at least twice the marginal expenditure, up to a small multiplicative error: $\dot{P}+\dot{C}\geq(2-\epsilon)\dot{S}$.
  2. (ii)
  3. (iii)
    The marginal expenditure is arbitrarily close to the target expenditure $e$, i.e., $|\dot{S}-e|<\epsilon$.
The condition in 1 stipulates that the market has $(M(n),\delta)$-recoverable structure for some $M(n)$ that is asymptotically much larger than $b(n)$. This lower bound on $M(n)$ ensures that the recoverable structure in $\bm{D}$ is substantially larger than the norm of the error matrix $\bm{E}$, which is $O(b(n))$ under 4.
Point (ii) states that it is possible to do this in a way that leaves all households’ welfare essentially unchanged; indeed, under the policy that we construct, the surplus gains are fully captured by the producers. Note that by 1, the welfare gain described in point (i) is essentially the maximum total surplus change that an omniscient authority could implement with the same amount of spending without reducing consumer surplus.
Our notion of $\epsilon$–robustness says that the surplus properties above are achieved ex post with high probability. A natural question is whether they are also achieved in expectation (since, in principle, realizations with very negative surplus could occur with low probability). In our setting, it turns out that all the analysis would be unaffected if we added good ex ante expected performance to the definition of robustness. [^18] This is because the surplus pass-throughs are supported on $[0,1]$ and all quantities in the proofs are bounded. Thus, convergence in probability is equivalent to convergence in $L^{1}$, and so our proofs extend to show close approximations to the omniscient benchmark in terms of ex ante surplus.
There are two main steps in the proof of 1. The first step is developed in Section 4.1 and uses the spectral decomposition (i.e., orthogonal diagonalization) of the Slutsky matrix to represent the pass-through of an intervention to prices, quantities and surpluses as a linear combination of orthogonal effects. Each of these orthogonal effects is associated with the projection of the intervention onto an eigenvector of $\bm{D}$. This approach highlights the potential to control the effects of an intervention by designing it in such a way that it projects exclusively onto subspaces spanned by eigenvectors of $\bm{D}$ that can be accurately identified.
In the second step, developed in Section 4.2, we apply a statistical method that allows us to accurately identify a subspace spanned by some eigenvectors of the Slutsky matrix from noisy observations. We then show that interventions that project exclusively onto these recoverable subspaces are robust and have the desirable welfare properties stated in 1.
4.1. Spectral decomposition and pass-throughs
4.1.1. Spectral decomposition of price and quantity effects
A key preliminary step is a spectral decomposition of the pass-through of an intervention to prices, quantities, and surpluses. In this section we present this decomposition from the perspective of an analyst who has perfect knowledge about the market state $(\bm{D},\bm{q}^{0})$.
Denoting by $\bm{U}$ the matrix whose $\ell^{\text{th}}$-column is the $\ell^{\text{th}}$ eigenvector $\bm{u}^{\ell}$ of $\bm{D}$, and by $\bm{\Lambda}$ the matrix whose non-diagonal elements are zero and whose $\ell^{\text{th}}$ diagonal element is $\lambda_{\ell}$, we have:
$\bm{D}=\bm{U}\bm{\Lambda}\bm{U}^{\mathsf{T}}.$
An intervention $\bm{\sigma}$ that subsidizes (or taxes) a single product will in general affect not only the prices and quantities of that product, but also those of other products, since in equilibrium these are all tied through strategic interactions. If we think of the eigenvectors $\bm{u}^{\ell}$ as representing bundles, then these bundles have the important property that an intervention $\bm{\sigma}\propto\bm{u}^{\ell}$ in the direction of such a bundle, will only affect the price $\bm{u}^{\ell}\cdot\bm{p}$ and quantity $\bm{u}^{\ell}\cdot\bm{q}$ of that bundle, leaving the prices and quantities of the bundles corresponding to the other eigenvectors unchanged. Generally, we can decompose $\bm{\sigma}=\sum_{\ell}(\bm{u}^{\ell}\cdot\bm{\sigma})\bm{u}^{\ell}$ into a combination of $n$ orthogonal interventions, each in the direction of an eigenvector. We can use this decomposition to solve the oligopoly game and get simple expressions for the pass-through of the intervention, in terms of the eigenvalues of the matrix.
Lemma 1.
Proof.
Thus, we can study the price and quantity pass-throughs of each of these $n$ interventions separably across eigenvectors. In particular, each unit of subsidy in direction $\bm{u}^{\ell}$ exclusively passes through to the price and quantities of bundle $\bm{u}^{\ell}$, and it does so with coefficients $-(1+|\lambda_{\ell}|)^{-1}$ and $|\lambda_{\ell}|(1+|\lambda_{\ell}|)^{-1}$, respectively.
Note that the magnitudes of the price and quantity pass-throughs in the different $\bm{u}^{\ell}$ are ordered according to their corresponding eigenvalues: The larger is $|\lambda_{\ell}|$, the less a given subsidy $\bm{u}^{\ell}\cdot\bm{\sigma}$ reduces prices, but the more it increases quantities. This asymmetry is the result of two opposing forces: On the one hand, the strategic interactions among firms imply that the equilibrium price $\bm{u}^{\ell}\cdot\bm{p}^{0}$ is less sensitive to the subsidy $\bm{u}^{\ell}\cdot\bm{\sigma}$ the larger is $|\lambda_{\ell}|$. On the other, the demand $\bm{u}^{\ell}\cdot\bm{q}^{0}$ is more sensitive to the price $\bm{u}^{\ell}\cdot\bm{p}^{0}$ the larger [^19] Indeed, it follows from (6) that $\bm{U}^{\mathsf{T}}\bm{q}^{0}=\bm{\Lambda}\bm{U}^{\mathsf{T}}\bm{p}^{0}$, so the slope of the demand $\bm{u}^{\ell}\cdot\bm{q}^{0}$ with respect to own price $\bm{u}^{\ell}\cdot\bm{p}^{0}$ is equal to $\lambda_{\ell}$. is $|\lambda_{\ell}|$; this is just a fact about the market’s demand function, rather than equilibrium pricing. 1 shows that the second effect dominates the first in the sense that the larger is $|\lambda_{\ell}|$, the more sensitive is the equilibrium quantity $\bm{u}^{\ell}\cdot\bm{q}^{0}$ to the subsidy $\bm{u}^{\ell}\cdot\bm{\sigma}$.
Lemma 2.
The pass-throughs to consumer, producer and total surpluses are:
Proof.
The effect of the intervention on consumer surplus is $\dot{C}=-\bm{q}^{0}\cdot\dot{\bm{p}}$. Multiplying this equation by $\bm{U}\bm{U}^{\mathsf{T}}$ gives $\dot{C}=-\bm{U}^{\mathsf{T}}\bm{q}^{0}\cdot\bm{U}^{\mathsf{T}}\dot{\bm{p}}$. Similarly, the effect of the intervention on producer surplus is $\dot{P}=\bm{q}^{0}\cdot(\dot{\bm{p}}+\bm{\sigma})+(\bm{p}-\bm{c})\cdot\dot{\bm{q}}=2\bm{q}^{0}\cdot\dot{\bm{q}}$. Multiplying the equation by $\bm{U}\bm{U}^{\mathsf{T}}$ we obtain the expression of $\dot{P}$ in the Lemma. The expression for $\dot{W}$ is obtained by aggregating $\dot{P}$, $\dot{C}$ and the expenditure of the intervention. ∎
2 shows that the effect of an intervention on consumer, producer or total surplus is a weighted sum of pass-throughs to each of the eigenvectors $\bm{u}^{\ell}$—with the weight being the corresponding bundle’s quantity.
We now use 2 to illustrate how—by setting $\bm{\sigma}$ equal to $\bm{u}^{1}$— the authority may be able to obtain the highest total surplus per-dollar spent possible subject to the constraint that the change in consumer surplus is not negative (recall 1). Because $\bm{u}^{1}$ is orthogonal to all the other $\bm{u}^{\ell}$ with $\ell\neq 1$, such intervention only changes the price and quantity of the bundle $\bm{u}^{1}$:
leading to an overall change in consumer and producer surplus equal to:
Theorem 1 focuses on the case of $M(n)\gg b(n)$, in which case $|\lambda_{1}|\rightarrow\infty$ as $n$ grows large. Consequently, the intervention has a vanishing effect (of magnitude $1/(1+|\lambda_{1}|)$) on equilibrium prices and consumer surplus. However, this intervention changes the quantity of bundle $\bm{u}^{1}$ by almost one, and the equilibrium quantity of each product $i$ changes proportionally to how much $i$ is represented in $\bm{u}^{1}$, i.e., $\bm{u}^{1}_{i}$. Furthermore, recoverable structure ensures that the projection of $\bm{u}^{1}$ on the status quo quantity vector is positive. This ensures that the intervention $\bm{\sigma}=\bm{u}^{1}$ has a non-zero level of expenditure, i.e., $\dot{S}=\bm{u}^{1}\cdot\bm{q}^{0}>0$. Overall, the intervention changes quantities so that the increase in total surplus is the maximum possible increase per dollar spent (as $n$ grows large), and these gains are captured entirely by the producers.
The next subsection explains that the property of recoverable structure allows us to estimate accurately objects like $\bm{u}^{1}$ from noisy estimates.
4.2. Recovering and using the subspace of dominant eigenvectors
[…]
4.2.1. The Davis–Kahan theorem
Recall that if $M(n)\gg b(n)$, then $(M(n),\delta)$-recoverable structure requires that the normalized Slutsky matrix $\bm{D}$ has eigenvalues that, in absolute value, are much larger than $b(n)$; henceforth, for brevity, we say such eigenvalues are “large.”
The key tool in our statistical exercise that leverages this assumption is the Davis–Kahan theorem. Under the hypothesis that some eigenvalues of $\bm{D}$ are large, this theorem guarantees that, despite the noise in $\bm{E}$, the large eigenvalues and associated eigenvectors of the observed matrix are a very good approximation of the true large eigenvalues. In other words, the noise in $\bm{E}$ cannot cause the large eigenvalues of $\bm{D}$ to be “mixed up” with the eigenvalues far away in the spectrum; see Figure 3 for an illustration. A bit more formally, this theorem has the following two central implications in our setting:
  1. (i)
  2. (ii)
    the eigenvectors of $\widehat{\bm{D}}$ associated with such eigenvalues are highly correlated with eigenvectors of $\bm{D}$ with similarly large eigenvalues, i.e., the eigenvectors of $\widehat{\bm{D}}$ associated with such eigenvalues can be written (up to a small error) as linear combinations of eigenvectors of $\bm{D}$ with similarly large eigenvalues.
This powerful result forms the core of our strategy to recover and use structure about the oligopoly demand in a way that is robust to noise. To facilitate the illustration, we make a stronger assumption on $\bm{D}$. The assumption is that the largest eigenvalue of $\bm{D}$ is sufficiently well-separated from all other eigenvalues by a “gap” much larger than $b(n)$. In that case, the Davis–Kahan Theorem has an even stronger implication: we can use $\widehat{\bm{D}}$ to recover a normalized eigenvector $\widehat{\bm{u}}^{1}$ that is almost perfectly positively correlated to the corresponding eigenvector $\bm{u}^{1}$. This stronger “gap” condition holds in Example 1 when $\gamma=0.3$ since the gap between the two largest eigenvalues is roughly $50$ and $b(n)$ is roughly $17$; see also Figure 1 as an illustration of the similarity between the true $\bm{u}^{1}$ (in panel B) and the estimated $\widehat{\bm{u}}^{1}$ (panel D).
To use this, note that as a consequence of 1 and 2, we have
while expenditure is
Let us use the recovered $\widehat{\bm{u}}^{1}$ to design an intervention with $\bm{\sigma}\propto\widehat{\bm{u}}^{1}$. The Davis–Kahan Theorem tells us that we can essentially take $\widehat{\bm{u}}^{1}$ to be its true counterpart $\bm{u}^{1}$ with a very small error, so from now on we will talk as if we know $\bm{u}^{1}$. Let us flip the sign of $\bm{\sigma}$ if necessary so that $(\bm{u}^{1}\cdot\bm{q}^{0})(\bm{u}^{1}\cdot\bm{\sigma})$ is positive. Then we can see from the equations above that $\dot{W}$ will in fact be very similar to $\dot{S}$, since $\lambda_{1}\gg b(n)$ is large and so $\frac{|\lambda_{1}|}{1+|\lambda_{1}|}\approx 1$. Moreover, if we know $\bm{u}^{1}\cdot\bm{q}^{0}$ and this differs from zero (which is a requirement of recoverable structure), we can scale the intervention to be of the size that we desire, and achieve $\dot{S}=1$.
The argument just described contains some wishful thinking, however. When we arranged the sign of $\bm{\sigma}$ so that $(\bm{u}^{1}\cdot\bm{q}^{0})(\bm{u}^{1}\cdot\bm{\sigma})$ is positive, we did not consider that we have only a noisy observation $\widehat{\bm{q}}^{0}$ of $\bm{q}^{0}$. So part of the challenge of the proof is to manage the observation error that makes $\widehat{\bm{q}}^{0}$ different from ${\bm{q}}^{0}$, and to show that we can obtain a correct estimate of the sign with probability tending to $1$ as $n\to\infty$. If we don’t manage to do this correctly, our intervention actually decreases efficiency with positive probability. This explains the need for the second part of Assumption 4 on $\bm{\varepsilon}$ in the signal.
However, assuming that noise is bounded is not enough: if $\bm{u}^{1}\cdot\bm{q}^{0}$ is very small, then there may be no hope for consistently recovering the true magnitude or sign of $\bm{u}^{1}\cdot\bm{q}^{0}$ from the signal $\widehat{\bm{u}}^{1}\cdot(\bm{q}^{0}+\bm{\varepsilon})$ even if the noise is well-behaved: the asymptotically small noise may still overwhelm a similarly decaying underlying mean $\bm{u}^{1}\cdot\bm{q}^{0}$. Such an unrecoverability would make it impossible to orient and scale our intervention appropriately. The definition of recoverable structure ensures that this problem does not occur by requiring that the projection of $\bm{q}^{0}$ onto eigenvectors with large eigenevalues is bounded away from zero.
The special case of the result that this discussion makes plausible is this: If the largest eigenvalue of $\bm{D}$ is well-separated from others and if $\bm{q}^{0}\cdot\bm{u}^{1}$ is not vanishingly small, then a subsidy profile that is proportional to $\bm{u}^{1}$ can, if it is suitably scaled, achieve all the properties of 1.
4.2.3. The more general result
The proof of the main result improves on this sketch in two ways. First, it does not impose that we need to recover exactly $\bm{u}^{1}$, but rather uses a much larger eigenspace of $\widehat{\bm{D}}$. Indeed, the general intervention projects $\widehat{\bm{q}}^{0}$ onto $\mathcal{L}(\bm{D},M(n))$, the eigenspace of all eigenvectors of $\bm{D}$ larger than $M(n)$. This makes it much easier for the analogue of $\bm{u}^{1}\cdot\bm{q}^{0}$ not to be too small, since the projection of $\bm{q}^{0}$ onto a larger eigenspace will have a larger norm. Second, the general proof dispenses with assuming that any eigenvalues are well-separated. Instead, it deals with any possible spectrum of $\bm{D}$ subject to our maintained assumptions. This is a considerable complication, because it is no longer generally possible to recover any true eigenvector $\bm{u}^{\ell}$ with any accuracy. We instead work directly with a recovered eigenspace $\mathcal{L}(\bm{D},M(n))$ that generalizes the span of $\widehat{\bm{u}}^{1}$ and show that while we know little about the individual true eigenvectors of $\bm{D}$ giving rise to this space, we can use the fact that all of them have large eigenvalues to generalize our argument for showing that (8) and (9) can be made very close and nonzero with a feasible intervention. This is where the arguments are more subtle and require going beyond standard applications of the Davis–Kahan theorem.
5. Illustration: Interventions based on noisy demand estimates
We present simulations that illustrate the central ideas behind 1. In particular, within the context of Example 1, we illustrate the effects of the intervention rule that taxes and subsidizes firms in the direction of the eigenvector associated with the largest eigenvalue of the observed Slutsky matrix; we do so by evaluating the effect of this intervention in many randomly generated market states.
5.1. Generating the true and observed market state
We construct the true and observed Slutsky matrices $\bm{D}$ and $\widehat{\bm{D}}$ as we described in Example 1. In particular, $n=300$, $\widehat{\bm{D}}=\bm{D}+\bm{E}$ and
$\bm{D}=(1-\gamma)\bm{D}_{\text{block}}+\gamma(-\bm{Z}\bm{Z}^{\prime}).$
Recall that $\bm{D}_{\text{block}}$ has three blocks of equal size, products in the same block are complements and products across block are substitutes. The parameter $\gamma\in(0,1)$ controls the size of the large eigenvalues of $\bm{D}$, with low values of $\gamma$ creating large eigenvalues by weighting the block matrix $\bm{D}$ whose eigenvalues grow linearly (and so much faster than $b(n)\sim n^{1/2}$).

### v2

@@ 5. Illustration: Interventions based on noisy demand estimates
The first matrix, $\bm{D}_{\text{block}}$, is a block matrix dividing the $300$ products into $3$ equally-sized blocks. The entries of $\bm{D}_{\text{block}}$ are based on a $3\times 3$ matrix $\bm{C}$ that governs the pattern of interactions within blocks $(C_{ii}$) and across them ($C_{ij}$ with $i\neq j$). In particular, [^24] The symbol $\otimes$ denotes the Kronecker product.
@@
Here $\bm{J}$ is the matrix of ones. Products within each block are complements, while products across blocks are substitutes. For instance, products in block 3 are highly substitutable with products in other blocks, whereas products in blocks 1 and 2 are only mildly substitutable with each other.
We can think of each block as a product category—for example, kitchen products, digital entertainment products, and sports-related products. Within categories, products are complements, but across the categories they are substitutes. For another example, consider three non-compatible operating systems accomplishing similar tasks; products within each system are complements, and products across them are substitutes.
The second term in eq. 12, a scaling of $-\bm{Z}\bm{Z}^{\prime}$, is a negative definite matrix that adds heterogeneity to the regular pattern of interactions in $\bm{D}_{\text{block}}$. In particular, we take $\bm{Z}$ to be an $n\times 10$ matrix with rows drawn uniformly from the unit sphere. [^25] Note that entry $(i,i)$ of $\bm{Z}\bm{Z}^{\prime}$ is the norm of the $i$th row of $Z$, which, by construction, is $1$.
@@
where each entry of the error matrix $\bm{E}$—except its diagonal entries, which are kept at zero—is drawn from $U[-1,1]$; this noise structure results in the upper bound $b(n)=n^{1/2}$ on $\|\bm{E}\|$.
We aim to illustrate basic statistical implications of $(M,\delta)$-recoverable structure that is strong enough relative to the noise $b(n)$—that is, with $M$ much larger than $b(n)$. In this example, the largest eigenvalue of $\bm{D}_{\text{block}}$ substantially exceeds $n^{1/2}$. Hence, if $\gamma$ is low, then the $\bm{D}$ of eq. 12 will also have some eigenvalues exceeding $n^{1/2}$. In contrast, if $\gamma$ is high, then the main contributor to $\bm{D}$ is $-\bm{Z}\bm{Z}^{\prime}$ and, in this case, the largest eigenvalues of $\bm{D}$ turn out to be no larger than $n^{1/2}$. We now show that these two cases, low vs.  high $\gamma$, have very different implications for inference.
@@
First we consider a low value of $\gamma=0.3$. In this case, the two largest eigenvalues of $\bm{D}$, in absolute value, are approximately $130$ and $80$, and these are considerably larger than $b(n)=\sqrt{300}\approx 17$, whereas the third largest eigenvalue is roughly $1.2$; hence, $\mathcal{L}(\bm{D},M)$ with $M>b(n)$ equal to, e.g., $n^{2/3}$, is the subspace spanned by eigenvectors $\bm{u}^{1}$ and $\bm{u}^{2}$. As we explained in Section 4.3, the Davis-Kahan Theorem tells us that we can use the noisy observation $\widehat{\bm{D}}$ to accurately estimate the eigenvectors in $\mathcal{L}(\bm{D},M)$. Here, for simplicity, we illustrate graphically how the first eigenvector of $\widehat{\bm{D}}$ is a good description of the first eigenvector of $\bm{D}$.
@@
5.1. Interventions
In the context of Example 2, we focus on the following intervention rule: Recover the eigenvector associated to the largest eigenvalue, in absolute terms, of the estimated Slutsky matrix $\widehat{\bm{D}}$. Intervene to subsidize firms in proportion to this eigenvector. This intervention aligns with the intervention rule behind our Theorem 1.262626In this example, for low values of $\gamma$ the largest eigenvalue of $\bm{D}$ is sufficiently well-separated from all other eigenvalues and, consequently, the authority can use the first eigenvector of $\widehat{\bm{D}}$ as a good approximation of $\bm{u}^{1}$ (see the illustrative discussion in Section 4.3.2).